# Introducing MeMo: A Multimodal Dataset for Memory Modelling in Multiparty Conversations

### v1

@@ Abstract
The quality of human social relationships is intricately linked to human memory processes, with memory serving as the foundation for the creation of social bonds. Since human memory is selective, differing recollections of the same events within a group can lead to misunderstandings and misalignments in what is perceived to be common ground in the group. Yet, conversational facilitation systems, aimed at advancing the quality of group interactions, usually focus on tracking users’ states within an individual session, ignoring what remains in each participant’s memory after the interaction. Conversational memory is the process by which humans encode, retain and retrieve verbal, non-verbal and contextual information from a conversation. Understanding conversational memory can be used as a source of information on the long-term development of social connections within a group. This paper introduces the MeMo corpus, the first conversational dataset annotated with participants’ memory retention reports, aimed at facilitating computational modelling of human conversational memory. The MeMo corpus includes 31 hours of small-group discussions on the topic of Covid-19, repeated over the term of 2 weeks. It integrates validated behavioural and perceptual measures, and includes audio, video, and multimodal annotations, offering a valuable resource for studying and modelling conversational memory and group dynamics. By introducing the MeMo corpus, presenting an analysis of its validity, and demonstrating its usefulness for future research, this paper aims to pave the way for future research in conversational memory modelling for intelligent system development.
@@
Suggesting potential topics of future research using the corpus. We describe some potential topics that can be explored with the use of the MeMo corpus in the areas specific to memory modelling, as well as wider fields of affective computing, intelligent system development, and behavioural research (see Section 8.3).
@@ 2 Motivation for creating MeMo
Existing conversation support systems have been shown to improve social interaction satisfaction, encourage equal participation and decrease social inhibitions [25, 26, 27]. A common method of achieving these results involves continuous tracking of users’ non-verbal and verbal signals [28]. These low-level signals are then used to infer a real-time measure of users’ participation (e.g.[26]) or more complex internal states such as attention (e.g. [29]), dominance (e.g. [30]) or social presence (e.g. [31]). Based on these predicted measures, a system then produces suggestions on how to enforce meeting structure and promote equal participation (e.g. [28], [26]). These predicted measures are real-time indicators of a participant’s immediate reactions or internal states, captured at specific timestamps or as cumulative measures from the start of the session. Although these measures can be used to represent users’ current state or trends within a session, they do not always represent the way the user will feel about the subject in subsequent interactions, since feelings triggered by an event within an interaction could be forgotten or completely changed over time, in retrospect to the event [32]. To sum up, user experience itself may not matter as much as the user’s memory of that experience in the context of long-term interaction and future desicions [33]. Consequently, for long-term interaction, a facilitation system needs to not only track the current state of the user but also understand the user’s memory of conversational experiences.
@@
Unlocking the potential of personal memories for user-modeling and personalization purposes in conversational settings requires predictive models based on real-world conversational data. In another context - that of media consumption - researchers have built such models to predict which images or videos are more likely to be encoded and retained by a human viewer based on media features and user characteristics [37, 38, 39, 40]. However, the conversational context is different from media consumption: unlike media consumption, conversational context involves continuous production and comprehension of verbal and non-verbal signals, involving different cognitive mechanisms [41] and producing qualitatively and quantitatively different memories [4]. Moreover, conversational memory has various context-specific factors at play that do not apply to media consumption tasks: conversation-specific verbal and non-verbal signals, the relational dynamics between conversational partners, and many more [3, 5, 6, 7, 8]. Therefore, the task of modelling how humans encode, retain and retrieve conversations remains unsolved.
@@ 3.1 P1: Maximising ecological validity
P1.3 Ensuring Representativeness of Participants. Lastly, it is important for a dataset to recruit a representative sample of participants from diverse demographics and backgrounds. It is, unfortunately, a common practice in datasets and experimental studies to mainly recruit university students and staff, biasing the data towards a demographic of highly educated English-speaking white young women [48, 49]. An alternative to university students is a more diverse demographic of participants recruited through specialised websites such as Amazon Mechanical Turk. While this method provides a more diverse demographic, participants recruited this way might be ’professional participants’ who go through a multitude of studies daily and are therefore biased in how they respond to the experiment questions [50]. In either case, study results can greatly depend on the demographics of its participants [51], and it is therefore important to report the demographics along with the dataset to understand the limitations of the data. So far, unfortunately, it is not a common practice and many datasets do not report the demographics of their participants (see section Section 4.5 for examples).
@@ 3.3 P3: Considering Context-sensitivity of Memory Processes
Specifically, conversational memory can be affected by communication skills [7], mood [56, 57], personality [58], values [59] and the relationship dynamics between participants [5, 60]. In addition, factors in relation to group perception should be measured such as group entitativity, cohesion and rapport. While they have not been investigated in relation to conversational memory the research shows that they can influence learning [61, 62], which is inherently related to memory.
@@ 4 Related work
To the best of our knowledge there are no datasets focusing on conversational memory. However, to contextualise MeMo, we describe existing datasets that are aimed to support computational modelling research on memory in one way or another. In addition, we describe most related behavioural studies on the topic of conversational memory. We compare the data designs using the criteria most relevant for the context of the MeMo corpus as shown in Table I. We describe each criterion in the following subsections.
@@ 4.1 Facilitated Modelling Perspective
When it comes to datasets for modelling memory processes, we distinguish between two different modelling perspectives that they facilitate: 1. Corpora supporting Situation-centered perspectives facilitate modelling how specific properties of a defined situation (e.g., exposure to a video) are expected to give rise to memory responses in members of some population (e.g., how specific video content is likely to be remembered by people in general). Datasets focusing on a situation-centered modelling perspective often attempt to capture a large range of distinct situations but typically have a small number of distinct individuals responding to them (e.g. [63]). In contrast, datasets with an 2. Individual-centeredperspective typically focus on more fine-grained modelling of variation in memory processes across specific individuals, possibly considering interactions with the situation (e.g., ways in which individuals behaviorally express when specific video content triggers a memory in them [64]). Datasets focusing on supporting an Individual-centered modelling perspective often contain only a relatively small number of distinct situations but a relatively large number of individuals responding to them. Note that datasets facilitating an Individual-centered perspective can often also support a Situation-cetered one, but not the other way around (because responses are typically aggregated from individual responses to the situation-level).
In MeMo, we aim to combine the two perspectives as much as the conversational context permits. From an individual-centered perspective, MeMo includes various perceptual and audio-visual measures from individual participants, providing resources to study how humans behave during memorable moments. From a situation-centered perspective, the data design allows for the investigation of the entire situation using audio-visual data from all group members, along with memory data aggregated across the group. This approach helps identify what makes a moment more memorable for a set of participants, abstracting from individual differences (for an example see Section 8.3).
@@ 4.2 Memory process
These processes are closely intertwined with each other. For example, any study of memory involves some measure of memory retention - whether or not a memory was preserved and for how long. Whether a moment has been retained or not cannot be completely measured, since some memories might have been retained but are not accessible at the moment of the measure [67]. Therefore, when it comes to retention, researchers usually focus on investigating memories available for retrieval at the moment of a memory test. Forgetting then refers to the moments that are not available at the moment of collecting the memory measure. Studies that focus on the retention process usually investigate the forgetting curve - collecting memory reports at several points in time and seeing how much information will be retained across the time [68]. Two datasets have collected such data in the context of media retention and forgetting [38, 69]. To our knowledge, there have been no studies investigating forgetting in free-flowing conversations.
@@
The MeMo dataset has been designed to study episodic memory two memory sub-processes. First, the corpus allows for studying when information has been encoded through the investigation of first-party temporal labels of events registered in participants’ free-recall reports (see Section 5.2). Second, the corpus is designed to study conversational retention, with memory reports collected immediately after interactions for short-term memory and after 3-4 days for longer-term memory (see more detail in Section 5.2). Although not its primary focus, MeMo could also be used for memory retrieval research by annotating past experiences mentioned in subsequent conversations.
@@ 6.2 Data processing
Memory data processing. We excluded the memory moments with invalid timestamps: if the reported start was further than the reported end of a memory event and if the reported timestamps were outside the duration of the recording. Since the multi-modal data processing included shifting the start and the end of the recordings, the memory temporal annotations needed to be adjusted to the new start and end of the recordings, to maintain the references to the originally tagged events.
@@ 7.3.1 Longitudinal completeness of questionnaire data
Although participants were asked to complete a questionnaire before and after each conversation session, there are some gaps in the data where participants could not complete questionnaires due to unforeseen circumstances. To facilitate the selection of the data for longitudinal analysis, we have computed the continuity of the questionnaire data in each group. Apart from two groups, every group had 50 or more participants with complete pre- and post-questionnaires in all the sessions. We share the questionnaire completeness scores for each session for an easier subset selection process in future research (see table in Appendix 5 for the full table and measure description).
@@ 8.1.1 Temporal dependency
To test the hypothesis, we compared the memorability index (percentage of participants that included the segment in their memory reports, see labels used in [89]) of moments that occured in the first 1/3 of the session, in the middle and in the end 1/3 of the session. Judging by an ANOVA followed by a paired t-test, the memorability index of the moments at the start and the end of the conversations is significantly higher than the ones in the middle (p0.005). This therefore seems to confirm the recency and primacy bias in relation to memorable moments occurrence. Therefore, the memory labels should not be treated as independent of temporal context within a session.
@@ 8.1.2 Session dependency
Another factor to consider in MeMo dataset relates to its’ longitudinal quality. Since each group participated in 3 consecutive sessions, starting as strangers and gradually getting to know each other, there might have been some evolution in their relationships and group perception. This matters for two reasons. First, a session-dependent variation would indicate the success of moderated sessions in creating a connection within the group ( G2) and show whether the MeMo corpus is representative of general societal trends. Second, this variability is important for computational modelling to determine whether the sessions can be treated as independent of each other. To investigate whether there is a consistent change in participants’ perceptions of each other and the group, we analysed self-reported ratings collected from participants after each session during the experiment.
@@ 8.2 Example: Group-level Conversational Memory Prediction
The first baseline for memorability prediction based on MeMo corpus has been explored and shown to be promising in [89]. Tsfasman et al. [89] used group gaze features to automatically predict group memorability levels.
@@ 8.3.1 Conversational memory encoding modelling
Task Relevance. The task of modelling the encoding of conversational moments using participants’ (non-)verbal behaviour at the time of the encoding could be relevant for several applications. For instance, in meeting facilitation systems, knowing which conversational segments are likely to be encoded by participants could allow systems to focus on moments that contribute most to participants’ mutual understanding, common ground and shared history [21, 91]. In addition, sharing this information with participants could highlight what one person found important but others overlooked, indicating that it might have to be revisited further in the discussion. In automatic meeting summarisation, identifying which segments are most likely to be encoded by participants could enable summaries that emphasise the information that was most important for each participant or augment users’ memory by focusing on the points that are less likely to be remembered [92, 93]. The benefit of using the moment of memory encoding for these applications is that it could potentially be done in real time, providing new conversational input.
@@ 8.3.2 Conversational memory retention modelling
Task Relevance. This task is particularly relevant for long-term facilitation systems or conversational agent applications. Specifically, since long-term user engagement with these systems remains a challenge [94, 95], there have been attempts to enhance conversational agents with shared memory models [96, 97, 98, 99]. However, there has not been a way to actually identify which memories are shared with the user and which ones a user would have forgotten. Memory retention modelling could help the issue, improving agents’ understanding of the user in the long term and adjusting dialogue strategies to the likelihood of conversational events being retained short-term or long-term. Facilitation systems could similarly benefit from understanding which conversational events are so important to the user that they are retained long-term. These memory models can be potentially modified with a dementia factor to support dementia patients, and, in general, be personalised for optimal user experience.
@@ 8.3.3 Perceived reason for retention modelling
8.3.4 Conversational memory retrieval modelling
Although not a primary focus of the dataset, the MeMo corpus offers data that could potentially be used for modelling spontaneous retrieval of conversational events in subsequent interactions. Since the MeMo corpus includes longitudinal data, modelling the spontaneous retrieval of memories from previous sessions is possible with the use of MeMo data. The research question would be what kind of memories are more likely to be revisited within the subsequent conversations and how they are used within the conversation. For this study to be possible, there needs to be an additional annotation that tags what utterances refer to an event in a previous interaction and an additional annotation of when that event occurred in the previous section. That annotation could then be linked to the encoding and retention data mentioned in the other modelling tasks above. This enables the study of modelling when to spontaneously retrieve information for the benefit of the group or individual.
@@ 10 Limitations
As with any dataset, the MeMo corpus has its limitations.
First, the MeMo corpus captures the memory of online video-call conversations, which could be different from the memory of face-to-face interaction, because of different conversational dynamics. For example, [100] reported shorter timing of turn-taking and more overlaps in an online setting compared to face-to-face conversations. The reason might originate in the fact that there is only one audio source in an online setting, while in a face-to-face setting every person provides a separate source of audio. This said the research on how humans remember information from online and face-to-face lectures shows similar recall quality has no differences between the settings [101].
Second, the topic of the conversations in the MeMo corpus is limited to the Covid-19 pandemic. At the time of the recording, this topic was naturally engaging since it noticeably affected most people’s lives. Humans tend to remember information that is personally relevant to them [102] and therefore with a different topic, the trends in memorable moments might have been different. In addition, a more goal-directed conversational context such as a work meeting could bring different results depending on the goals and role in the conversation.
Third, the longitudinal character of the MeMo corpus is limited to three sessions over two weeks. In addition, the presence of a trained moderator might add a certain constraint and hierarchy to the group organisation. Having moderators in the interaction could also add a confounding factor of different moderation styles and therefore varying discussion structure.
Lastly, the temporal annotation of memories assumes that there is one specific moment in the conversation. Therefore, there might be some events that were included in the memory but are not captured by the annotation. In addition, identifying the segment in the recording is a cognitively taxing task, which might make the segment annotation less precise, as the participant might already be tired from the earlier tasks.

### v2

@@ Abstract
Conversational memory is the process by which humans encode, retain and retrieve verbal, non-verbal and contextual information from a conversation. Since human memory is selective, differing recollections of the same events can lead to misunderstandings and misalignments within a group. Yet, conversational facilitation systems, aimed at advancing the quality of group interactions, usually focus on tracking users’ states within an individual session, ignoring what remains in each participant’s memory after the interaction. Understanding conversational memory can be used as a source of information on the long-term development of social connections within a group. This paper introduces the MeMo corpus, the first conversational dataset annotated with participants’ memory retention reports, aimed at facilitating computational modelling of human conversational memory. The MeMo corpus includes 31 hours of small-group discussions on Covid-19, repeated 3 times over the term of 2 weeks. It integrates validated behavioural and perceptual measures, audio, video, and multimodal annotations, offering a valuable resource for studying and modelling conversational memory and group dynamics. By introducing the MeMo corpus, analysing its validity, and demonstrating its usefulness for future research, this paper aims to pave the way for future research in conversational memory modelling for intelligent system development.
@@
Suggesting potential topics of future research using the corpus. We describe potential modelling tasks that can be explored with the use of the MeMo corpus (see Section 8.3).
@@ 2 Motivation for creating MeMo
Existing conversation support systems have been shown to improve social interaction satisfaction, encourage equal participation and decrease social inhibitions [25, 26, 27]. A common method of achieving these results involves continuous tracking of users’ non-verbal and verbal signals [28]. These low-level signals are then used to infer a real-time measure of users’ participation (e.g.[26]) or more complex internal states such as attention (e.g. [29]), dominance (e.g. [30]) or social presence (e.g. [31]). Based on these predicted measures, a system then produces suggestions on how to enforce meeting structure and promote equal participation (e.g. [28], [26]). These predicted measures are real-time indicators of a participant’s immediate reactions or internal states, captured at specific timestamps or as cumulative measures from the start of the session. Although these measures can be used to represent users’ current state or trends within a session, they do not always represent the way the user will feel about the subject in subsequent interactions, since feelings triggered by an event within an interaction could be forgotten or completely changed over time, in retrospect to the event [32]. To sum up, user experience itself may not matter as much as the user’s memory of that experience in the context of long-term interaction and future decisions [33]. Consequently, for long-term interaction, a facilitation system needs to not only track the current state of the user but also understand the user’s memory of conversational experiences.
@@
Unlocking the potential of personal memories for user-modelling and personalisation purposes in conversational settings requires predictive models based on real-world conversational data. In another context - that of media consumption - researchers have built such models to predict which images or videos are more likely to be encoded and retained by a human viewer based on media features and user characteristics [37, 38, 39, 40]. However, the conversational context is different from media consumption: unlike media consumption, conversational context involves continuous production and comprehension of verbal and non-verbal signals, involving different cognitive mechanisms [41] and producing qualitatively and quantitatively different memories [4]. Moreover, conversational memory has various context-specific factors at play that do not apply to media consumption tasks: conversation-specific verbal and non-verbal signals, the relational dynamics between conversational partners, and many more [3, 5, 6, 7, 8]. Therefore, the task of modelling how humans encode, retain and retrieve conversations remains unsolved.
@@ 3.1 P1: Maximising ecological validity
P1.3 Ensuring Representativeness of Participants. Lastly, it is important for a dataset to recruit a representative sample of participants from diverse demographics and backgrounds. It is, unfortunately, a common practice in datasets and experimental studies to mainly recruit university students and staff, biasing the data towards a demographic of highly educated English-speaking white young women [48, 49]. An alternative to university students is a more diverse demographic of participants recruited through specialised websites such as Amazon Mechanical Turk. While this method provides a more diverse demographic, participants recruited this way might be ’professional participants’ who go through a multitude of studies daily and are therefore biased in how they respond to the experiment questions [50]. In either case, study results can greatly depend on the demographics of its participants [51], and it is, therefore, important to report the demographics along with the dataset to understand the limitations of the data. So far, unfortunately, it is not a common practice and many datasets do not report the demographics of their participants (see section Section 4.5 for examples).
@@ 3.3 P3: Considering Context-sensitivity of Memory Processes
Specifically, conversational memory can be affected by communication skills [7], mood [56, 57], personality [58], values [59] and the relationship dynamics between participants [5, 60]. In addition, factors concerning group perception should be measured such as group entitativity, cohesion and rapport. While they have not been investigated in relation to conversational memory the research shows that they can influence learning [61, 62], which is inherently related to memory.
@@ 4 Related work
To the best of our knowledge, there are no datasets focusing on conversational memory. However, to contextualise MeMo, we describe existing datasets that are aimed to support computational modelling research on memory in one way or another. In addition, we describe most related behavioural studies on the topic of conversational memory. We compare the data designs using the criteria most relevant for the context of the MeMo corpus as shown in Table I. We describe each criterion in the following subsections.
@@ 4.1 Facilitated Modelling Perspective
When it comes to datasets for modelling memory processes, we distinguish between two different modelling perspectives that they facilitate: 1. Corpora supporting Situation-centered perspectives facilitate modelling how specific properties of a defined situation (e.g., exposure to a video) are expected to give rise to memory responses in members of some population (e.g., how specific video content is likely to be remembered by people in general). Datasets focusing on a situation-centred modelling perspective often attempt to capture a large range of distinct situations but typically have a small number of distinct individuals responding to them (e.g. [63]). In contrast, datasets with an 2. Individual-centeredperspective typically focus on more fine-grained modelling of variation in memory processes across specific individuals, possibly considering interactions with the situation (e.g., ways in which individuals behaviorally express when specific video content triggers a memory in them [64]). Datasets focusing on supporting an Individual-centered modelling perspective often contain only a relatively small number of distinct situations but a relatively large number of individuals responding to them. Note that datasets facilitating an Individual-centered perspective can often also support a Situation-centred one, but not the other way around (because responses are typically aggregated from individual responses to the situation level).
In MeMo, we aim to combine the two perspectives as much as the conversational context permits. From an individual-centred perspective, MeMo includes various perceptual and audio-visual measures from individual participants, providing resources to study how humans behave during memorable moments. From a situation-centered perspective, the data design allows for the investigation of the entire situation using audio-visual data from all group members, along with memory data aggregated across the group. This approach helps identify what makes a moment more memorable for a set of participants, abstracting from individual differences (for an example see Section 8.3).
@@ 4.2 Memory process
These processes are closely intertwined with each other. For example, any study of memory involves some measure of memory retention - whether or not a memory was preserved and for how long. Whether a moment has been retained or not cannot be completely measured, since some memories might have been retained but are not accessible at the moment of the measure [67]. Therefore, when it comes to retention, researchers usually focus on investigating memories available for retrieval at the moment of a memory test. Forgetting then refers to the moments that are not available at the moment of collecting the memory measure. Studies that focus on the retention process usually investigate the forgetting curve - collecting memory reports at several points in time and seeing how much information will be retained across the time [68]. Two datasets have collected such data in the context of media retention and forgetting [38, 69]. To our knowledge, there have been no studies investigating retention or forgetting in free-flowing conversations.
@@
The MeMo dataset has been designed to study and model two sub-processes of human episodic memory. First, the corpus allows for studying when information has been encoded through the investigation of first-party temporal labels of events registered in participants’ free-recall reports (see Section 5.2). Second, the corpus is designed to study conversational retention, with memory reports collected immediately after interactions for short-term memory and after 3-4 days for longer-term memory (see more detail in Section 5.2; also notice the limitations for this task in Section 10). In principle, MeMo could also be used for modelling memory retrieval processing during conversations (e.g., in terms of how people use memories for social bonding purposes during interactions [1]). However, the annotations provided with the current release do explicitly support this task (see Section 10 for a discussion).
@@ 6.2 Data processing
Memory data processing. We excluded the memory moments with invalid timestamps: if the reported start was further than the reported end of a memory event and if the reported timestamps were outside the duration of the recording. Since the multi-modal data processing included shifting the start and the end of the recordings, the encoded event annotations needed to be adjusted to the new start and end of the recordings, to maintain the references to the originally tagged events.
@@ 7.3.1 Longitudinal completeness of questionnaire data
Although participants were asked to complete a questionnaire before and after each conversation session, there are some gaps in the data where participants could not complete questionnaires due to unforeseen circumstances. To facilitate the selection of the data for longitudinal analysis, we have computed the continuity of the questionnaire data in each group. Apart from the two groups, every group had 50 or more participants with complete pre- and post-questionnaires in all the sessions. We share the questionnaire completeness scores for each session for an easier subset selection process in future research (see table in Appendix 5 for the full table and measure description).
@@ 8.1.1 Temporal dependency
To test the hypothesis, we compared the memorability index (percentage of participants that included the segment in their memory reports, see labels used in [89]) of moments that occurred in the first 1/3 of the session, in the middle and in the end 1/3 of the session. Judging by an ANOVA followed by a paired t-test, the memorability index of the moments at the start and the end of the conversations is significantly higher than the ones in the middle (p0.005). This therefore seems to confirm the recency and primacy bias in relation to memorable moments occurrence. Therefore, the memory labels should not be treated as independent of temporal context within a session.
@@ 8.1.2 Session dependency
Another factor to consider in the MeMo dataset relates to its’ longitudinal quality. Since each group participated in 3 consecutive sessions, starting as strangers and gradually getting to know each other, there might have been some evolution in their relationships and group perception. This matters for two reasons. First, a session-dependent variation would indicate the success of moderated sessions in creating a connection within the group ( G2) and show whether the MeMo corpus is representative of general societal trends. Second, this variability is important for computational modelling to determine whether the sessions can be treated as independent of each other. To investigate whether there is a consistent change in participants’ perceptions of each other and the group, we analysed self-reported ratings collected from participants after each session during the experiment.
@@ 8.2 Example: Group-level Conversational Memory Prediction
The first baseline for memorability prediction based on the MeMo corpus has been explored and shown to be promising in [89]. Tsfasman et al. [89] used group gaze features to automatically predict group memorability levels.
@@ 8.3.1 Conversational memory encoding modelling
Task Relevance. Modelling how conversational moments are encoded through participants’ (non-)verbal behaviour could be useful for various applications. In meeting facilitation systems, identifying which segments are likely encoded by participants can help focus on moments that enhance mutual understanding and shared history [21, 91]. Sharing this information might highlight points one person found important but others missed, prompting further discussion. In automatic meeting summarisation, recognising which segments are likely encoded can produce summaries that emphasise key information or enhance memory by focusing on less memorable points [92, 93].
@@ 8.3.2 Conversational memory retention modelling
Task Relevance. This task is relevant for long-term facilitation systems and conversational agents. Since long-term user engagement remains a challenge [94, 95], efforts have been made to enhance agents with shared memory models [96, 97, 98, 99]. However, there is no way to identify which memories are shared or forgotten by the user. Memory retention modelling could improve agents’ long-term understanding of users and adapt dialogue strategies based on the likelihood of events being retained. Facilitation systems could also benefit from knowing which events are retained long-term. These models could be personalised for better user experience or modified to support dementia patients.
@@ 10 Limitations
As with any dataset, the MeMo corpus has its limitations. In the following, we will highlight several key aspects relevant to the intended use of the corpus.
Validity of Memory Annotations
MeMo contains two types of memory annotations: Free-Recall Self-reports (FRS), and Encoded Event Annotations (EEAs). While we argue that jointly these provide valid resources for modelling both the encoding and retention processes of participants, there are some caveats we want to highlight.
First, our corpus assumes that each FRS corresponds to a single EEA, i.e., a remembered moment in the conversation corresponds to a single segment in the multimodal signal. However, this might not always be true, for example, when multiple similar events are reported as a single memory (with more than one EEA included in one FRS).
Secondly, it is likely that there are events that participants remembered, but failed to either report or annotate. Our protocol required participants to identify multiple segments in the recording at a fine granularity, which is a cognitively taxing task. This might impact the comprehensiveness and accuracy of the resulting annotations.
Finally, our protocol asked participants to report FRS from previous sessions (see Section 5.2.4). We have argued that this data forms a viable source for modelling retention. However, we cannot rule out that participants could access memories from previous sessions but failed to report them (i.e., we do not know if they forgot about a moment or simply failed to produce a matching FRS).
Comprehensiveness of Memory Annotations
While we believe that FRS and EEAs capture properties relevant for modelling retention and encoding, the current annotations do not facilitate work on the functional use of memories in conversations (i.e., how people use memories to convince others or for bonding purposes [1]). However, it is likely that MeMo captures a substantial amount of such instances, and we intend to investigate their presence in future work (i.e., to provide additional annotations). Note that even though our protocol for obtaining FRS explicitly asks participants to ”recall” memory, this is merely a methodological necessity (i.e., it is not possible to study memory content without prompting for it); it does not facilitate modelling retrieval processes as such (i.e., the conditions under which content is dynamically accessed or accessible).
Choice of Conversational Setting
First, the MeMo corpus was recorded online. Therefore, it is specific to the memory of online video-call conversations, which could be different from the memory of face-to-face interaction, because of different conversational dynamics [100]. This said, the research on how humans remember information from online and face-to-face lectures shows no differences in recall quality between these settings [101].
Second, the topic of the conversations in the MeMo corpus is limited to the Covid-19 pandemic. At the time of the recording, this topic was naturally engaging since it noticeably affected most people’s lives. Humans tend to remember information that is personally relevant to them [102] and therefore with a different topic, the trends in memorable moments might have been different.
Additionally, having a trained moderator may create a sense of hierarchy in the group as well as introduce different moderation styles, potentially affecting the discussion structure and group dynamics. A different setting might lead to different results, depending on the environment, goals, and roles in the conversation.